OpenStudy (anonymous):

Find the dimensions of the rectangle of largest area that has its base on the x-axis and its other two vertices above the x-axis and lying on the parabola. y=5-x^2

OpenStudy (dumbcow):

length of rectange is represented as 2x, where x is x_value of point on parabola width is represented as y, y_value of point on parabola Area = 2x*y = 2x(5-x^2) = 10x -2x^3 maximize Area by finding x value where derivative is zero dA/dx = 10 -6x^2 = 0 --> x = sqrt(5/3) optimal dimensions: length = 2sqrt(5/3) width = 10/3

OpenStudy (anonymous):

thank you